OpenStudy (anonymous):

In C[-pi, pi] , consider the set {1/Sqrt[2], cos(2x), sin(2x)} and define the inner product as =1/Pi \[\int_{a}^{b} f(x) g(x) dx\]

OpenStudy (anonymous):

a) find ||1/sqrt[2]||

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

but \[\frac{1}{\sqrt{2}}\] is a number right, so you are treating it as a constant function

OpenStudy (amistre64):

reminds me of the fourier stuff ive been looking at

OpenStudy (amistre64):

orthogonal functions and what not ....

OpenStudy (anonymous):

i take it you mean inner product is \[<f,g>=\frac{1}{\pi}\int_{-\pi}^{\pi}f(x)g(x)dx\]

OpenStudy (anonymous):

not the inner product equals \(\frac{1}{\pi}\) and not the integral from some a to some b

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

ok so replace both \(f\) and \(g\) by \(\frac{1}{\sqrt{2}}\)

OpenStudy (anonymous):

because \(||f||=<f,f>\)

OpenStudy (anonymous):

\[\int_{-pi}^{pi} 1/2 dx\]

OpenStudy (anonymous):

i made a mistake, norm is root of that

OpenStudy (anonymous):

yea, that is right \[\frac{1}{\pi}\int_{-\pi}^{\pi}\frac{1}{2}dx\]

OpenStudy (anonymous):

i think you get 1

OpenStudy (anonymous):

actually i am sure you get 1

OpenStudy (anonymous):

\[\frac{1}{\pi}\times 2\pi\times \frac{1}{2}=1\]

OpenStudy (anonymous):

and \(\sqrt{1}=1\)

OpenStudy (anonymous):

nice , thanks

OpenStudy (anonymous):

yw, my pleasure

OpenStudy (anonymous):

each of these has norm 1 i think